# Joint Channel, Data and Radar Parameter Estimation for AFDM Systems in Doubly-Dispersive Channels

### v2

@@ IV-A Reference SotA Method: SBL via Expectation Maximization
$=(\mathbf{F}_{M}\!\otimes\!\mathbf{I}_{K})\!\cdot\!\sigma^{2}_{w}\mathbf{I}_{N}\!\cdot\!(\mathbf{F}_{M}\!\otimes\!\mathbf{I}_{K})^{\mathsf{H}}\!=\!\sigma^{2}_{w}\Big{[}\mathbf{F}_{M}\mathbf{F}_{M}^{\mathsf{H}}\!\otimes\!\mathbf{I}_{K}\mathbf{I}_{K}^{\mathsf{H}}\Big{]}$
$=\sigma^{2}_{w}\Big{[}\mathbf{I}_{M}\otimes\mathbf{I}_{K}\mathbf{I}_{K}^{\mathsf{H}}\Big{]}=\sigma^{2}_{w}\mathbf{I}_{N}\in\mathbb{C}^{N\times N}.$ (56)
$=(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)\!\cdot\!\sigma^{2}_{w}\mathbf{I}_{N}\!\cdot\!(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)^{\mathsf{H}}\!=\!\sigma^{2}_{w}\mathbf{I}_{N}\in\mathbb{C}^{N\times N}.$
@@ IV-B Proposed Method: PDA-based Message Passing Estimator
As implied by the notation, the parameter vector $\bm{\theta}^{(i)}$ must be updated iteratively, which in similarity to the SBL approach of [30] can be accomplished via an EM algorithm. Unlike the aforementioned SotA method, however, the estimates of $\mathbf{h}$ in the proposed scheme are obtained via a message-passing algorithm, which is described in the sequel.
@@ IV-B4 Parameter Update via EM
To this end, we follow [47] and first convert $p_{\mathbf{y}|\bm{\theta}}(\mathbf{y}|\bm{\theta})$ to a tractable form by utilizing latent variables, marginalizing over an arbitrary PDF $\hat{p}_{\mathbf{{h}}}(\mathbf{h})$ and using the log-likelihood function $\text{ln}\big{[}p_{\mathbf{y}|\bm{\theta}}(\mathbf{y}|\bm{\theta})\big{]}$, reformulated so as to highlight the differential entropy and Kullback-Leibler (KL) divergence terms $\psi(\mathbf{h})$ and $\mathrm{D}_{\text{KL}}\big{(}\hat{p}_{\mathbf{{h}}}(\mathbf{h})||p_{\mathbf{h}|\mathbf{y},\bm{\theta}}(\mathbf{h}|\mathbf{y},\bm{\theta})\big{)}$, respectively, namely
@@
A complete and compact description of the proposed PDA-based RPE scheme is given in the form of a pseudocode in Algorithm (3). Upon convergence of the above procedure, the estimate of the sparse vector $\mathbf{h}$ is known. Finally, the $(i,j)$ indices of $\mathbf{h}$ can be used to obtain the corresponding delays and Dopplers, and hence the corresponding ranges and velocities of the objects in the vicinity.
IV-C Performance Assessment
In order to evaluate the performance of the proposed PDA-based RPE algorithm, we consider the monostatic scenario depicted in Fig. 1. Under such a scenario, and in light of the results of Figure 4 – as well as related literature, $e.g.$[8] – which demonstrate that OFDM is outperformed by OTFS and AFDM in doubly-dispersive channels, we consider hereafter only OTFS and AFDM systems.
Our first set of results, shown in 20, are obtained for a single LoS reflected signal, from a single target, positioned at $15$ m and traveling at a velocity of $151$ km/h towards the BS, with the remark that since both the SotA SBL and the proposed PDA RPE algorithms are based on the sparse model described by the set of equations (54), the estimation algorithm themselves have no hard restrictions on the number of paths or targets to be estimated, as long as the grid modeled by the dictionary matrix $\mathbf{E}$ in equation (54b) is fine enough to ensure a sufficient sparsity of the vector $\mathbf{y}$ in equation (54a).
@@ IV-C Performance Assessment
The performance metric adopted is the normalized root mean square error (RMSE), averaged over all targets and realizations, which is defined as
@@
Finally, Figure 20 also includes plots of the resolution limit [12] and the MF bound [48], obtained by executing Algorithm 3 for a single iteration, initialized with the true radar parameter values. The results clearly indicate that the proposed PDA method outperforms the SotA SBL approach both for range and velocity estimation, both under OTFS and AFDM waveforms, with the proposed method reaching the MF bound in the case of range estimates. It is also found, moreover, that AFDM yields slightly better performance than OTFS, and that all algorithms reach the resolution bound under a sufficiently large signal-to-noise ratio (SNR), indicating that both methods converge absolutely.
(a) Range Estimation Performance (FD system, single target).
(b) Velocity Estimation Performance (FD system, single target).
Figure 5: RMSE versus SNR  [^20] While many formulations for the SNR exist in literature, including radar-centric metrics as used in [8, 12], we use the typical definition hinging on $h_{p}$ as used in [20].performance of SotA SBL and proposed PDA monostatic RPE schemes over OTFS and AFDM waveforms in a FD (interference-free) scenario with a single LoS signal from a single target.
@@
(a) Range Estimation Performance (SIC system, multiple targets).
(b) Velocity Estimation Performance (SIC system, multiple targets).
Figure 7: RMSE versus SNR performance of SotA SBL and proposed PDA monostatic RPE schemes over AFDM waveform, subjected to self-interference removed via de-chirping as in [20], in a scenario with multiple targets.
@@
Having established the overall advantage of AFDM over OTFS, and of the proposed PDA-based RPE algorithm over the SBL-based SotA alternative, we finally confront the two critical issues left unaddressed in the performance assessment so far, namely, the impact of self-interference and a larger number of targets to be estimated. To that end, we offer in Figure 7 plots equivalent to those shown in Figure 20, but this time exclusively for AFDM systems, and in scenarios with multiple targets and subjected to self-interference, mitigated via the technique proposed in [20].
@@ V Conclusion
We offered two new ISAC-enabling contributions for the emerging AFDM waveform, namely, a PBiGaBP-based JCDE algorithm, which enables channel estimation with low pilot overhead, and a PDA-based RPE algorithm, which enables the BS to act as a monostatic radar, estimating the delay and Dopper parameters of multiple paths simultaneously. Using a generalized channel model under which OFDM, OTFS and AFDM can be directly compared, it was shown that the proposed algorithms outperform SotA alternatives and that AFDM is advantageous over OTFS and OFDM both for communications and sensing functionalities in doubly-dispersive channels.
@@
Figure 8: Diagram of Proposed PBiGaBP-based JCDE in Algorithm 1.
Figure 9: Diagram of Proposed PDA-based RPE Algorithm 3.

### v4

@@ IV-A Reference SotA Method: SBL via Expectation Maximization
$=(\mathbf{F}_{M}\!\otimes\!\mathbf{I}_{K})\!\cdot\!N_{0}\mathbf{I}_{N}\!\cdot\!(\mathbf{F}_{M}\!\otimes\!\mathbf{I}_{K})^{\mathsf{H}}\!=\!N_{0}\Big{[}\mathbf{F}_{M}\mathbf{F}_{M}^{\mathsf{H}}\!\otimes\!\mathbf{I}_{K}\mathbf{I}_{K}^{\mathsf{H}}\Big{]}$
$=N_{0}\Big{[}\mathbf{I}_{M}\otimes\mathbf{I}_{K}\mathbf{I}_{K}^{\mathsf{H}}\Big{]}=N_{0}\mathbf{I}_{N}\in\mathbb{C}^{N\times N}.$ (56)
@@
$=(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)\!\cdot\!N_{0}\mathbf{I}_{N}\!\cdot\!(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)^{\mathsf{H}}\!=\!N_{0}\mathbf{I}_{N}\in\mathbb{C}^{N\times N}.$
@@ IV-B Proposed Method: PDA-based Message Passing Estimator
As implied by the notation, the parameter vector $\bm{\theta}^{(i)}$ must be updated iteratively, which in similarity to the SBL approach of [31] can be accomplished via the EM algorithm. Unlike the aforementioned SotA method, however, the estimates of $\mathbf{h}$ in the proposed scheme are obtained via a message-passing algorithm, which is described in the sequel.
@@ IV-B4 Parameter Update via EM
To this end, we follow [63] and first convert $p_{\mathbf{y}|\bm{\theta}}(\mathbf{y}|\bm{\theta})$ to a tractable form by utilizing latent variables, marginalizing over an arbitrary PDF $\hat{p}_{\mathbf{{h}}}(\mathbf{h})$ and using the log-likelihood function $\text{ln}\big{[}p_{\mathbf{y}|\bm{\theta}}(\mathbf{y}|\bm{\theta})\big{]}$, reformulated to highlight the differential entropy and Kullback-Leibler (KL) divergence terms $\psi(\mathbf{h})$ and $\mathrm{D}_{\text{KL}}\big{(}\hat{p}_{\mathbf{{h}}}(\mathbf{h})||p_{\mathbf{h}|\mathbf{y},\bm{\theta}}(\mathbf{h}|\mathbf{y},\bm{\theta})\big{)}$, respectively, namely
@@
A complete and compact description of the proposed PDA-based RPE scheme is given in the form of a pseudocode in Algorithm (3). Upon convergence of the above procedure, the estimate of the sparse vector $\mathbf{h}$ is known. Finally, the $(i,j)$ indices of $\mathbf{h}$ can be used to obtain the corresponding delay and Doppler shifts, and hence the corresponding ranges and velocities of the objects in the vicinity.
Remark 2: Due to the nature of the RPE formulation, a state evolution analysis [64, 65] is no longer possible since the equivalent observation matrix is not independent and identically distributed (i.i.d.) sub-Gaussian distributed and/or right-unitary invariant with a finite system size. In addition, in order to handle the quite strong correlation amongst the elements of the equivalent observation matrix, the proposed RPE algorithm is designed based on the PDA framework [66], which while robust against such observation correlations, does not guarantee Bayesian optimality in the large-system limit. If the algorithm was designed via a Bayes-optimal technique such as generalized approximate message passing (GAMP) [67] or expectation propagation (EP) [68], the Onsager correction term will behave in an unstable manner and the estimation performance will deteriorate significantly. Given the above, it is difficult to perform a theoretical convergence analysis; hence, we have verified the number of iterations required for convergence via a thorough numerical analysis as shown below (Figure 7).
IV-C Complexity Analysis
The computational complexity for both the proposed PDA-based RPE scheme detailed in Algorithm 3, and the SotA SBL-based RPE scheme given in Algorithm 2, are dominated by the matrix inversion on line 2 of both algorithms, required at every iteration of the corresponding algorithms, albeit of different sizes in each.
While the SotA method has a computational complexity of order $\mathcal{O}((K_{\tau}D_{\nu})^{3})$, the complexity order of the proposed method is $\mathcal{O}(N^{3})$. In most cases, since $N\approx K_{\tau}\cdot D_{\nu}$, the overall computational complexity is very similar. It is also noteworthy that this translates to the proposed PDA-based RPE technique being almost completely independent of the grid size based on the values of $K_{\tau}$ and $D_{\nu}$, which means that much larger grids (even over multiple iterations as in [15]) for high accuracy can be used with the proposed method as opposed to the SotA which incurs a heavy computational burden due to the matrix inversion dependent on $K_{\tau}\cdot D_{\nu}$.
Finally, we emphasize that in addition to the performance gain and the similar complexity of the proposed method, we also show that the proposed method converges much faster, as shown in Figure 7.
IV-D Performance Assessment
In order to evaluate the performance of the proposed PDA-based RPE algorithm, we consider the monostatic scenario depicted in Figure 4. Under such a scenario, and in light of the results of Figure 3 – as well as related literature, $e.g.$[11]
(a) Range Estimation Performance (FD system, single target).
(b) Velocity Estimation Performance (FD system, single target).
Figure 5: RMSE versus SNR  [^24] While many formulations for the SNR exist in literature, including radar-centric metrics as used in [11, 15], we use the typical definition hinging on $h_{p}$ as used in [23], which already incorporates the effect of the radar cross-section parameter $\sigma_{\text{rcs}}=1$ denoting that no power is lost due to the reflection.performance of SotA SBL and proposed PDA monostatic RPE schemes over OTFS and AFDM waveforms in a FD (interference-free) scenario with a single LoS signal from a single target.
– which demonstrate that OFDM is outperformed by OTFS and AFDM in doubly-dispersive channels, we consider hereafter only OTFS and AFDM systems.
Our first set of results, shown in Figure 24, are obtained for a single LoS reflected signal, from a single target, positioned at $15$ m and traveling at a velocity of $151$ km/h towards the BS, with the remark that since both the SotA SBL and the proposed PDA RPE algorithms are based on the sparse model described by the set of equations (54), the estimation algorithm themselves have no hard restrictions on the number of paths or targets to be estimated, as long as the grid modeled by the dictionary matrix $\mathbf{E}$ in equation (54b) is fine enough to ensure a sufficient sparsity of the vector $\mathbf{y}$ in equation (54a).
@@ IV-D Performance Assessment
The performance metric adopted is the normalized RMSE, averaged over all targets and realizations, which is defined as
@@
Finally, Figure 24 also includes plots of the resolution limit [^25] In this case, the resolution limit defines how fine the search grid is; $i.e.,$ it measures the estimation accuracy for the range and velocity if the estimate for $\!\hat{\;\mathbf{h}}$, obtained as the output of the SotA and proposed algorithms, is perfect. [15] and the MF bound [69], obtained by executing Algorithm 3 for a single iteration, initialized with the true radar parameter values. The results clearly indicate that the proposed PDA method outperforms the SotA SBL approach both for range and velocity estimation, both under OTFS and AFDM waveforms, with the proposed method reaching the MF bound in the case of range estimates. It is also found, moreover, that AFDM yields slightly better performance than OTFS, and that all algorithms reach the resolution bound under a sufficiently large SNR, indicating that both methods converge absolutely.
@@
(a) Range Estimation Performance (SIC-enabled system, multiple targets).
(b) Velocity Estimation Performance (SIC-enabled system, multiple targets).
Figure 7: RMSE versus SNR performance of SotA SBL and proposed PDA monostatic RPE schemes over AFDM waveform, executed using only the single pilot and null-guard interval enabling SIC via a simple low-complexity analog dechirping and filtering[23], in a scenario with multiple targets.
@@
Having established the overall advantage of AFDM over OTFS in FD scenarios, and of the proposed PDA-based RPE algorithm over the SBL-based SotA alternative, we finally confront the two critical issues left unaddressed in the performance assessment so far, namely, the impact of self-interference and a larger number of targets to be estimated. To that end, we offer in Figure 7 plots equivalent to those shown in Figure 24, but this time exclusively for SIC-enabled AFDM systems, and in scenarios with multiple targets using only the single pilot and null-guard interval for RPE. Notice that this only requires minimal modification to equation (54) such that it becomes $\mathbf{y}_{p}=\mathbf{E}_{p}\cdot\mathbf{h}+\tilde{\mathbf{w}_{p}}\in\mathbb{C}^{B\times 1}$ with $\mathbf{y}_{p}\in\mathbb{C}^{B\times 1}$ and $\mathbf{E}_{p}\in\mathbb{C}^{B\times K_{\tau}D_{\nu}}$ being the parts of the received signal $\mathbf{y}\in\mathbb{C}^{N\times 1}$ and dictionary matrix $\mathbf{E}\in\mathbb{C}^{N\times K_{\tau}D_{\nu}}$, respectively, corresponding to the single pilot and the null-guard interval $\mathbf{x}_{p}\in\mathbb{C}^{B\times 1}$ of the original transmit symbol vector $\mathbf{x}\in\mathbb{C}^{N\times 1}$. Consequently, both the SotA and proposed algorithms apply to monostatic RPE, such that the inherent self-interference from the LoS path can be cancelled via a simple low-complexity analog dechirping and filtering[23] as opposed to the costly solutions required for OFDM and OTFS systems.
@@ V Conclusion
We offered two new ISAC-enabling contributions for the emerging AFDM waveform, namely, a PBiGaBP-based JCDE algorithm, which enables channel estimation with low pilot overhead, and a PDA-based RPE algorithm, which enables the BS to act as a monostatic radar, estimating the delay and Doppler parameters of multiple paths simultaneously. Using a generalized channel model under which OFDM, OTFS and AFDM can be directly compared, it was shown that the proposed algorithms outperform SotA alternatives and that AFDM is advantageous over OTFS and OFDM both for communications and sensing functionalities in doubly-dispersive channels.
@@
Figure 8: Diagram of the Proposed PBiGaBP-based JCDE in Algorithm 1.
Figure 9: Diagram of the Proposed PDA-based RPE in Algorithm 3.
[Uncaptioned image] Kuranage Roche Rayan Ranasinghe (Graduate Student Member, IEEE) obtained B.Sc. degrees in Electrical and Computer Engineering, as well as Robotics and Intelligent Systems (with a minor in Physics) from Constructor University (formerly Jacobs University Bremen) in 2023/2024, where he is currently pursuing a Ph.D. in Electrical Engineering. Within the fields of wireless communications and signal processing, his research interests encompass integrated sensing, communications and computing (ISCC), compressed sensing, Bayesian inference, belief propagation and optimization theory.
[Uncaptioned image] Hyeon Seok Rou (Graduate Student Member, IEEE) has received the Ph.D. degree in Electrical Engineering in 2024 from Constructor University, Bremen, Germany, and the B.Sc. degree in Electrical and Computer Engineering (ECE) in 2021 from Jacobs University Bremen, Germany. He has received the Korea Institute of Science and Technology Europe Research Scholarship Award from The Korean Scientists and Engineers Association in the FRG (Verein Koreanischer Naturwissenschaftler und Ingenieure in der BRD e.V.) in 2022, and he was a visiting researcher at the Intelligent Communications Lab, Korea Advanced Institute of Science and Technology (KAIST) in 2023. His research interests include integrated sensing and communications (ISAC), signal processing in doubly-dispersive channels, high-mobility communication systems, multi-dimensional modulation, next-generation metasurfaces, B5G/6G V2X communication technologies, and quantum computing.
[Uncaptioned image] Giuseppe Thadeu Freitas de Abreu (Senior Member, IEEE) received the B.Eng. degree in electrical engineering and the specialization Latu Sensu degree in telecommunications engineering from the Universidade Federal da Bahia (UFBa), Salvador, Bahia, Brazil in 1996 and 1997, respectively, and the M.Eng. and D.Eng. degrees in physics, electrical, and computer engineering from Yokohama National University, Japan, in March 2001 and March 2004, respectively. He was a postdoctoral fellow and later an adjunct professor (docent) in statistical signal processing and communications theory at the Department of Electrical and Information Engineering, University of Oulu, Finland from 2004 to 2006 and from 2006 to 2011, respectively. Since 2011, he has been a professor of electrical engineering at Jacobs University, Bremen, Germany. From April 2015 to August 2018, he simultaneously held a full professorship at the Department of Computer and Electrical Engineering, Ritsumeikan University, Japan. His research interests include communications and signal processing, including communications theory, estimation theory, statistical modeling, wireless localization, cognitive radio, wireless security, MIMO systems, ultrawideband and millimeter wave communications, full-duplex and cognitive radio, compressive sensing, energy harvesting networks, random networks, connected vehicles networks, and many other topics. He received the Uenohara Award at Tokyo University in 2000 for his master’s thesis. He has been a co-recipient of the Best Paper Award at several international conferences. He was awarded prestigious JSPS, Heiwa Nakajima, and NICT Fellowships in 2010, 2013, and 2015, respectively. He served as an associate editor for the IEEE Transactions on Wireless Communications from 2009 to 2014 and the IEEE Transactions on Communications from 2014 to 2017; and as an executive editor for IEEE Transactions on Wireless Communications from 2017 to 2021. He is currently serving as an editor to the IEEE Signal Processing Letters and the IEEE Communications Letters.
[Uncaptioned image] Takumi Takahashi (Member, IEEE) received the B.E., M.E., and Ph.D. degrees in communication engineering from Osaka University, Osaka, Japan, in 2016, 2017, and 2019, respectively. From 2018 to 2019, he was a Visiting Researcher at the Centre for Wireless Communications, University of Oulu, Finland. In 2019, he joined the Graduate School of Engineering, Osaka University, as an Assistant Professor. His current research interests include Bayesian inference, belief propagation, signal processing, and wireless communications. He received the 80th Best Paper Award from IEICE and the 2019 and 2023 Best Paper Awards from IEICE Communication Society. He was certified as an Exemplary Reviewer of IEEE Wireless Communications Letters in 2023.
[Uncaptioned image] Kenta Ito (Graduate Student Member, IEEE) received the B.E. degree in electrical engineering from Toyama University, Japan, in 2019, and the M.E. degree in communication engineering from Osaka University, Japan, in 2021. He is currently pursuing the Ph.D. degree at the Graduate School of Engineering, Osaka University. His research interests include belief propagation, compressed sensing, signal processing, and wireless communications.